OpenStudy (anonymous):

squared root negative 15 cubed/2 cubed times the variable x + -(40.56)= 20 - 7 cubed/ 2 squared

OpenStudy (anonymous):

its 20 mixed number -7 to the power of 3/ 2 squared

OpenStudy (anonymous):

it's not clear what you've written there... to make your expressions clearer, use notation like: (15)^3 / (2)^3 if necessary, use parenthesis to indicate numerators and denominators as well: ( (15)^3 )/( (2)^3 )

OpenStudy (anonymous):

its very clear

OpenStudy (anonymous):

cubed stands for 3rd power squared is to the second 2 power

OpenStudy (anonymous):

the square root is the bracket foundation for the rest of the problem so right everything else within the square root symbol

OpenStudy (anonymous):

or just use the equation editor for great justice

OpenStudy (anonymous):

sadly idk how it works

OpenStudy (anonymous):

ok write a square root symbol with 15 to the third power as the numerator and 2 cubed as the denominator times the variable x + -40.56= 20 mixed number 17 numerator/ division sign 2 to the second power i cant make it any clearer than that

OpenStudy (anonymous):

I mean -7 to the third power division sign 2 to the second power

OpenStudy (anonymous):

squared root negative 15 cubed/2 cubed times the variable x + -(40.56)= 20 - 7 cubed/ 2 squared sqrt((-15^3/2^3) x + - 40.56 = 20 - (7^3/2^2)? <-- the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to use square u need to type ^2 in the equation table =)

OpenStudy (anonymous):

(-) square root is an imaginary number. sqrt((-15^3/2^3) x